OpenStudy (anonymous):

The formula K equals C plus two hundred seventy-three shows how to estimate the temperature K measured in degrees Kelvin (°K) when you know the temperature C measured in degrees Celsius (°C). If room temperature is 295°K, what is room temperature in °C?

OpenStudy (anonymous):

\[K=C+273\]Replace the value of K by 295 and solve it for C\[295=C+273\]

OpenStudy (anonymous):

295 = c + 273 c = 295 - 273 c = 22

OpenStudy (anonymous):

postive or negative 22

OpenStudy (anonymous):

positive

OpenStudy (anonymous):

okay thank you guys so much

OpenStudy (anonymous):

your welcome